OpenStudy (anonymous):

Evaluate e^In when x=3 and y=2

OpenStudy (zarkon):

there are no x's or y's in your expression

OpenStudy (anonymous):

i mean e^xIny

OpenStudy (zarkon):

\[e^{x\ln(y)}\] this?

OpenStudy (anonymous):

yes

OpenStudy (zarkon):

\[e^{x\ln(y)}=e^{\ln(y^x)}=y^x\] now replace y by 2 and x by 3

OpenStudy (anonymous):

so the answer wil be 8

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

thank you very much

OpenStudy (zarkon):

np

OpenStudy (anonymous):

what about this logx(1/4)=-2 find x

OpenStudy (zarkon):

is the 1/4 inside the log?

OpenStudy (zarkon):

\[\log\left(x\frac{1}{4}\right)=-2\]

OpenStudy (anonymous):

no the logx is seperate fro 1/4

OpenStudy (zarkon):

\[\log\left(x\right)\frac{1}{4}=-2\]

OpenStudy (anonymous):

yes

OpenStudy (zarkon):

is this log base 10?

OpenStudy (anonymous):

what do you mean

OpenStudy (zarkon):

\[\log_{a}(b)\] what is \(a\) in this case?

OpenStudy (anonymous):

it just x and nothing else